ABSORPTION SPECTROCOPY OF Σ0_{0} → Π1_{0} BAND IN Rg·ND_{3} (Rg = Ne, Ar, Kr)

\underline{DMITRY G. MELNIK}, MYKOLA BATAIEV, TERRY A. MILLER, The Ohio State University, Dept. of Chemistry, Laser Spectroscopy Facility, 120 W. 18th Avenue, Columbus, Ohio 43210; and FRANK C. DE LUCIA, The Ohio State University, Dept. of Physics, Microwave Laboratory, 174 W. 18th Avenue, Columbus OH 43210.

Previously, we have reported the Σ0_{0} → Π1_{0} vibrational-tunneling-rotational (VTR) transition in the Ar·ND_{3} complex.^{a} Here we report the first observation of the corresponding transition in in the Ne·ND_{3} and Kr·ND_{3} complexes using a pulsed jet FASSST spectrometer operating in the 195-298 GHz region. The inversion components of the VTR band corresponding to different nuclear spin states, are observed for all complexes Rg·ND_{3} (Rg = Ne, Ar, Kr). The experimental data were fit to a pseudo-diatomic Hamiltonian,^{b} yielding molecular constants, including the value of the inversion splitting, Δ, in the complexes. A semiquantitative model of the (moderate) quenching of the inversion splitting in the complex is proposed.

aD.Melnik, S.Gopalakrishnan, T.A.Miller, and F.C.De Lucia, 54^{th} International Symposium on Molecular Spectroscopy

bY. Xu and A. R. W. McKellar, Mol.Physics 88, 859 (1996)